A proton enters a constant magnetic field of magnitude 0.0550 T and traverses a semicircle of radius 1.28 mm before leaving t will rate, thanks!
Answer #1

11.264 7103 Solution -27 6.67X10 Mp= kg -19 1.6 X/0 с qp = B - 00550T r= 1.28 mm ra mo LS When Protion initial clinchen ů Per